Please help me determine whether the equation is always, sometimes, or never true
6x-12+2x=3+8x-15

Found 2 solutions by ikleyn, MathTherapy:
Answer by ikleyn(52787)

.
6x - 12 + 2x = 3 + 8x - 15
--------------------------

This equation is always true.

It is an identity.

To get the conclusion, simplify this equation as you usually do.
It means: collect the terms with x; collect the constant terms etc.

In this way you will get

(6x + 2x - 8x) + (-12 -3 + 15) = 0, or

0*x + 0 = 0, or

0 = 0.

This equation is always true, so the original equation is.


Answer by MathTherapy(10552)


Please help me determine whether the equation is always, sometimes, or never true
6x-12+2x=3+8x-15
Always TRUE, since when solved, results in: 0 = 0.